# On quasi-Albanese morphisms for log canonical Calabi-Yau pairs

### v1

@@ Abstract.
We study the quasi-Albanese morphisms for log canonical Calabi-Yau pairs.
@@ 1. Introduction
Let $X$ be a complex projective variety with canonical singularities and $K_{X}$ torsion. Kawamata proved that its Albanese morphism $\alpha:X\to A$ is an étale fiber bundle [Kaw85, Theorem 8.3]. Inspired by the works [Iit76, Kaw81, Zha05, Fuj24, BFPT24], we study quasi-Albanese morphisms for (the open part of) log canonical Calabi-Yau pairs.
Before stating the main theorem, we recall the definition of logarithmic Kodaira dimension and a theorem by Kawamata and Fujino. Let $X^{0}$ be a smooth quasi-projective variety. Let $X$ be a compactification of $X^{0}$ such that $D_{X}:=X\setminus X^{0}$ is a simple normal crossing divisor. The logarithmic Kodaira dimension $\overline{\kappa}(X^{0})$ of $X^{0}$ is defined as $\kappa(K_{X}+D_{X})$. By Iitaka [Iit77], $\overline{\kappa}(X^{0})$ is independent of the choice of such $X$.
Theorem 1.
[Kaw81, Thoerem 28], see also [Fuj24, Theorem 1.2] If $\overline{\kappa}(X^{0})=0$, then the quasi-Albanese morphism of $X^{0}$ is dominant and has irreducible general fibers.
The following is our main theorem.
Theorem 2.
Let $(X,D_{X})$ be a log smooth pair with $D_{X}$ reduced and $K_{X}+D_{X}\sim_{\mathbb{Q}}0$. Let $X^{0}=X\setminus D_{X}$ and let $\alpha:X^{0}\to G$ be its quasi-Albanese morphism. Then
@@
    $\alpha$ is surjective in codimension one: for any prime divisor $P\subset G$, there is a prime divisor $Q\subset X^{0}$ dominating $P$,
@@
    $\alpha$ is flat in codimension one: for any prime divisor $Q\subset X^{0}$, if $Q$ do not dominate $G$, then $\alpha(Q)$ is of codimension one,
@@
    $\alpha$ is semistable in codimension one: for any prime divisor $P\subset G$, the pair $(X,\alpha^{*}P)$ is log canonical near the generic point of $P$.
Theorem 2 is a consequence of the following theorem, whose proof depends on Theorem 1 and the techniques from the minimal model program.
Theorem 3.
Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced and $K_{X}+D_{X}\sim_{\mathbb{Q}}0$. Set $q(X):=\dim H^{1}(O_{X})$ and $q(X,D_{X}):=\dim H^{0}(\Omega_{X^{0}}^{1}(\log D_{X}|_{X^{0}}))$, where $X^{0}$ is the log smooth locus of $(X,D_{X})$. Let $A$ be the Albanese variety of $X$ and let $G$ be the quasi-Albanese variety of $X\setminus D_{X}$. Then there are
  • a $\mathbb{Q}$-factorial dlt pair $(W,D_{W})$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$,
  • $(q(X,D_{X})-q(X))$’s line bundles $L_{i}\in\mathrm{Pic}^{0}(A)$, and
  • a commutative diagram,
@@
    where $\mathbb{P}_{A}:=\mathbb{P}_{A}(O_{A}\oplus L_{1}^{-1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}^{-1})$ is a compactification of $G$, $H=\mathbb{P}_{A}\setminus G$ is a simple normal crossing divisor, $g$ is a fibration, and $p$ is the natural projection,
such that the following holds.
  1. (1)
    $p\circ g$ is the Albanese morphism of $W$, $h$ is the quasi-Albanese morphism of $W\setminus D_{W}$.
  2. (2)
    The $g$-vertical part of $D_{W}$ coincides with $g^{-1}H$.
  3. (3)
    The canonical bundle formula for $g:(W,D_{W})\to\mathbb{P}_{A}$ has the form
    $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H),$
    and the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable.
As a corollary, if $q(X)=0$ and $q(X,D_{X})=\dim X$, then $(X,D_{X})$ is crepant birational to $(\mathbb{P}^{n},(\prod_{i}z_{i}=0))$. Another corollary is that if $q(X,D_{X})-q(X)\geq 2$, then $D_{X}$ is connected.
In the previous Theorem, the pair $(W,D_{W})$ is closely related to $(X,D_{X})$. For example, they have the same orbifold fundamental group, and the same dual complex (up to a PL-homeomorphism) (cf. [dFKX17, Proposition 11]). On the other hand, the fibration $g:(W,D_{W})\to\mathbb{P}_{A}$ has nice properties, which may reduce some questions about $(X,D_{X})$ to the case $q(X,D_{X})=0$ by induction on dimension. We give an example in the last section.
Acknowledgments
The author thanks professors Zhan Li, Lei Zhang, and Junchao Shentu for useful discussions. The author is supported by USTC.
2. Preparations
We work over the field of complex numbers. For the singularities of pairs and the minimal model program, we refer to the book [KM98].
A projective surjective morphism between normal varieties with connected fibers is called a fibration. Let $f:X\to Y$ be a fibration between normal varieties. Let $D$ be an effective $\mathbb{Q}$-divisor on $X$. We can write $D=D^{h}+D^{v}$, where $D^{v}$ does not dominate $Y$, and each component of $D^{h}$ dominates $Y$. $D^{v}$ (resp. $D^{h}$) is called the vertical part (resp. horizontal) part of $D$. $D$ is called vertical (resp. horizontal) if $D=D^{v}$ (resp. $D=D^{h}$). Assume that $D$ is vertical, then $D$ is called exceptional over $Y$ if $\operatorname{codim}f(D)\geq 2$, and very exceptional over $Y$ if for any prime divisor $P\subset Y$, there is a prime divisor $Q\subset X$ such that $Q$ dominates $P$ but $Q\not\subset D$.
For a birational map $\pi:X\dashrightarrow Y$ between normal varieties, we set $\operatorname{Ex}(\pi):=\sum P$, where the sum is taken over all prime divisors on $X$ which are exceptional over $Y$. A birational map $\pi:X\dashrightarrow Y$ is called a birational contraction if $\operatorname{Ex}(\pi^{-1})=0$.
For a projective variety $X$, we set $q(X):=\dim H^{1}(O_{X})$, and for a projective lc pair $(X,D_{X})$, we set $q(X,D_{X}):=\dim H^{0}(\Omega_{X^{0}}^{1}(\log\lfloor D_{X}|_{X^{0}}\rfloor))$, where $X^{0}$ is the log smooth locus of $(X,D_{X})$.
2.1. Differential forms on lc pairs
Theorem 4 ([GKKP11]).
Let $(X,D_{X})$ be an lc pair. Let $X^{0}$ be its log smooth locus and let $\mathrm{nonklt}(X,D_{X})$ be its non-klt locus. Let $\pi:Y\to X$ be a log resolution of $(X,D_{X})$ and let $D_{Y}$ be the codimension one part of $\pi^{-1}\mathrm{nonklt}(X,D_{X})$. Then $\pi_{*}\Omega_{Y}^{k}(\log D_{Y})$ is reflexive and $H^{0}(Y,\Omega_{Y}^{k}(\log D_{Y}))\cong H^{0}(X^{0},\Omega_{X^{0}}^{k}(\log\lfloor D_{X}|_{X^{0}}\rfloor))$ for all $k$.
Proposition 5.
Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced. Let $\pi:Y\to X$ be a log resolution of $(X,D_{X})$ and let $D_{Y}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. Then
  1. (1)
    $q(X)=q(Y)$,
  2. (2)
    $q(X,D_{X})=q(Y,D_{Y})$,
  3. (3)
    $\kappa(K_{X}+D)=\kappa(K_{Y}+D_{Y})$ and $\nu(K_{X}+D)=\nu(K_{Y}+D_{Y})$,
where $\kappa$ and $\nu$ are Kodaira dimension and numerical Kodaira dimension respectively.
@@ 2.1. Differential forms on lc pairs
Since klt varieties have rational singularities, we have (1). Since $(X,D_{X})$ is lc, there is an effective exceptional divisor $E$ on $Y$ such that $K_{Y}+D_{Y}=\pi^{*}(K_{X}+D_{X})+E$, so we have (3).
Now we prove (2). Note that $D_{Y}\geq\pi^{-1}D_{X}$ and $D_{X}=\mathrm{nonklt}(X,D_{X})$. By Theorem 4, to show (2), it suffices to show that the injective homomorphism
$\iota:H^{0}(\Omega_{Y}^{1}(\log\pi^{-1}D_{X}))\to H^{0}(\Omega_{Y}^{1}(\log D_{Y}))$
is an isomorphism. Write $\pi^{-1}D_{X}=\sum_{i}P_{i}$ and $D_{Y}-\pi^{-1}D_{X}=\sum_{j}E_{j}$ as sum of prime divisors. By the exact sequence
one checks that $\iota$ is an isomorphism iff $\sum_{i}a_{i}c_{1}(O_{Y}(P_{i}))+\sum_{j}b_{j}c_{1}(O_{Y}(E_{j}))=0$ implies that all $b_{j}$’s are zero. Which is indeed the case since all $E_{j}$’s are exceptional and $X$ is $\mathbb{Q}$-factorial.∎
Lemma 6.
Let $(X,D_{X})$ and $(Y,D_{Y})$ be two projective $\mathbb{Q}$-factorial dlt pairs with reduced $D_{X}$ and $D_{Y}$. Assume that there is a birational contraction $\pi:X\dashrightarrow Y$ such that $D_{X}=\pi^{-1}_{*}D_{Y}+\operatorname{Ex}(\pi)$ then $q(X,D_{X})=q(Y,D_{Y})$.
Proof.
Let $X\xleftarrow{p}W\xrightarrow{q}Y$ be a common log resolution. Then $q^{-1}_{*}D_{Y}+\operatorname{Ex}(q)=p^{-1}_{*}D_{X}+\operatorname{Ex}(p)$. Hence, the Lemma follows from the previous proposition. ∎
2.2. Quasi-Albanese morphism
Theorem 7 (Quasi-Albanese morphism [Iit76], see also [Fuj24]).

### v2

@@ Abstract.
In this note, we study quasi-Albanese morphisms for log canonical Calabi-Yau pairs and obtain several structural results. As an application, we prove a characterization of toric pairs.
@@ 1. Introduction
Let $X$ be a smooth complex projective Calabi-Yau variety. By the Bogomolov-Beauville-Yau decomposition, the Albanese morphism of $X$ is isotrivial. Kawamata [kawamata1985minimal, Theorem 8.3] extended this isotriviality to varieties with canonical singularities, and Ambro [Ambro2005moduli, Theorem 4.8] to klt pairs. For log canonical Calabi-Yau pairs, Bernasconi, Filipazzi, Patakfalvi and Tsakanikas [bernasconi2024strong, Theorem 1.5] proved that the Albanese morphism is locally stable in the sense of Kollár [Kollar-Families, Definition-Theorem 4.7], but may not be isotrivial. In the setting of log pairs, it is more natural to consider logarithmic differential forms and quasi-Albanese morphisms by Iitaka [iitaka1976logarithmic], which capture finer structural information than the Albanese morphisms.
In this note, we study quasi-Albanese morphisms for log canonical Calabi-Yau pairs. Our results are intended as a step toward a log canonical Bogomolov–Beauville–Yau decomposition.
We briefly recall the relevant definitions. A quasi-abelian variety $G$ is a quasi-projective commutative algebraic group fitting into an exact sequence
$0\to(\mathbb{C}^{*})^{d}\to G\to A\to 0,$
where $A$ is an abelian variety and $(\mathbb{C}^{*})^{d}$ is an algebraic torus; we refer to $A$ as the compact part of $G$. Let $X$ be a quasi-projective variety. The quasi-Albanese morphism $\alpha:X\to G$ is a morphism to a quasi-abelian variety characterized by the following universal property: any morphism from $X$ to a quasi-abelian variety factors uniquely through $\alpha$. The variety $G$ is called the quasi-Albanese variety of $X$.
We recall the following classical result by Kawamata.
Theorem 1.1 ([kawamata1981characterization, Theorem 28], see also [fujino2024quasi, Theorem 1.2]).
Let $(X,D_{X})$ be a projective log smooth pair with $D_{X}$ reduced and $\kappa(K_{X}+D_{X})=0$. Then the quasi-Albanese morphism $\alpha:X\setminus D_{X}\to G$ is dominant and has irreducible general fibers.
We now turn to the singular setting. For a log canonical pair $(X,D_{X})$, we denote by $\operatorname{nklt}(X,D_{X})$ its non-klt locus. We strengthen Kawamata’ Theorem as follows.
Theorem 1.2.
Let $(X,D_{X})$ be a projective log canonical pair with $\kappa(K_{X}+D_{X})=0$. Then the quasi-Albanese morphism $\alpha:X\setminus\operatorname{nklt}(X,D_{X})\to G$ is dominant and has irreducible general fibers.
The next theorem further analyzes the quasi-Albanese morphism $\alpha$ in the Calabi-Yau case.
Theorem 1.3.
Let $(X,D_{X})$ be a projective log canonical pair with $K_{X}+D_{X}\sim_{\mathbb{Q}}0$. Let $\alpha:X\setminus\operatorname{nklt}(X,D_{X})\to G$ be the quasi-Albanese morphism. Then the following holds.
@@
    $\alpha$ is surjective in codimension one: for any prime divisor $P\subset G$, there is a prime divisor $Q\subset X^{0}$ dominating $P$.
@@
    $\alpha:(X\setminus\operatorname{nklt}(X,D_{X}),D_{X}|_{X\setminus\operatorname{nklt}(X,D_{X})})\to G$ is locally stable. In particular, $\alpha$ is flat and every fiber is reduced with semi-log canonical singularities.
Let $G$ be a quasi-abelian variety and let $A$ be its compact part. Then there are $d:=\dim G-\dim A$ line bundles $L_{i}\in\text{Pic}^{0}(A)$ such that $G$ is the fiber product over $A$ of the total spaces of the $L_{i}$ with their zero sections removed. Consequently, $G$ admits the following natural compactification:
$G\subset\mathbb{P}_{A}:=\mathbb{P}_{A}(O_{A}\oplus L_{1}^{-1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}^{-1}).$
Note that $H:=\mathbb{P}_{A}\setminus G$ is a simple normal crossing divisor, and the log smooth pair $(\mathbb{P}_{A},H)$ is Calabi-Yau: $K_{\mathbb{P}_{A}}+H\sim 0$. We refer to $(\mathbb{P}_{A},H)$ as the canonical compactification of $G$.
Let $(X,D_{X})$ be a log canonical pair. Via the quasi-Albanese morphism $\alpha:X\setminus\operatorname{nklt}(X,D_{X})\to G$, we aim to extract structural information about the pair $(X,D_{X})$ itself, rather than merely about the open subset $X\setminus\operatorname{nklt}(X,D_{X})$. To this end, we consider the canonical compactification $(\mathbb{P}_{A},H)$ of $G$ and the induced rational map $X\dashrightarrow\mathbb{P}_{A}$. In general, however, this rational map need not be a morphism (see Example 2.5). The following theorem asserts that for a Calabi-Yau pair $(X,D_{X})$, one can always find another Calabi-Yau pair $(W,D_{W})$, crepant birational to $(X,D_{X})$, such that the induced map $W\dashrightarrow\mathbb{P}_{A}$ is a morphism.
Theorem 1.4.
Let $(X,D_{X})$ be a projective log canonical pair with $K_{X}+D_{X}\sim_{\mathbb{Q}}0$. Let $\alpha:X\setminus\operatorname{nklt}(X,D_{X})\to G$ be the quasi-Albanese morphism. Let $(\mathbb{P}_{A},H)$ be the canonical compactification of $G$. Then there are
  • a projective $\mathbb{Q}$-factorial dlt pair $(W,D_{W})$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$, and
  • a crepant birational contraction $\mu:(W,D_{W})\dashrightarrow(X,D_{X})$ such that $D_{W}=\mu^{-1}_{*}D_{X}+\sum P$, where the sum runs over all prime divisors on $W$ that are exceptional over $X$,
satisfying the following properties:
@@
    The induced rational map $g:W\dashrightarrow\mathbb{P}_{A}$ is a morphism, which is surjective and has connected fibers.
  2. (2)
    The canonical bundle formula for $g:(W,D_{W})\to\mathbb{P}_{A}$ takes the form
    $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H),$
    and the restriction $(W,D_{W})|_{g^{-1}(G)}\to G$ is locally stable.
  3. (3)
    The support of $g^{*}H$ is contained in the non-klt locus $\operatorname{nklt}(W,D_{W})$ and the induced morphism
    $\beta:W\setminus\operatorname{nklt}(W,D_{W})\to G$
    is the quasi-Albanese morphism of $W\setminus\operatorname{nklt}(W,D_{W})$.
As a byproduct of our study, we find a connection between the Shokurov complexity of log canonical pairs and their quasi-Albanese varieties. For a log canonical pair $(X,D_{X})$, its complexity $c(X,D_{X})$ is defined as
$c(X,D_{X}):=\dim X+\operatorname{rank}\langle\lfloor D_{X}\rfloor\rangle-|\lfloor D_{X}\rfloor|,$
where
  • $\langle\lfloor D_{X}\rfloor\rangle$ is the subgroup generated by the components of $\lfloor D_{X}\rfloor$ in the group of Weil divisors modulo algebraic equivalence, and
  • $|\lfloor D_{X}\rfloor|$ is the sum of the coefficients of $\lfloor D_{X}\rfloor$.
For simplicity, we adopt a definition slightly different from that in [Brown2018GeomToric]. In that work, Brown, McKernan, Svaldi, and Zong proved a conjecture of Shokurov [Shokurov2000CompleSurf] characterizing toric varieties via complexities. See also [Prokhorov2001Toric, yao2013thesis, Nakayama2017Shokurov, enwright2024log] for related results.
In this note, we show the following.
Lemma 1.5.
Let $(X,D_{X})$ be a projective log canonical pair. Let $G$ be the quasi-Albanese variety of $X\setminus\operatorname{nklt}(X,D_{X})$, and let $A$ be its compact part. Then
$c(X,D_{X})=\dim X-(\dim G-\dim A).$
Theorem 1.6.
Let $(X,D_{X})$ be a projective log canonical pair with $\kappa(K_{X}+D_{X})=0$. Then, for any reduced divisor $0\leq\Delta_{X}\leq\lfloor D_{X}\rfloor$, we have
$c(X,\Delta_{X})\geq 0.$
Assume that the equality holds for some $\Delta_{X}$. Then $X$ is rational. Assume further that $K_{X}+D_{X}\sim_{\mathbb{Q}}0$, then the following holds.
  1. (1)
    $\Delta_{X}=\lfloor D_{X}\rfloor=D_{X}$.
  2. (2)
    The quasi-Albanese morphism
    $\alpha:X\setminus\operatorname{nklt}(X,D_{X})\to(\mathbb{C}^{*})^{n}$
    is birational and isomorphic in codimension one.
  3. (3)
    There is a crepant birational map
    $(X,D_{X})\dashrightarrow(\mathbb{P}^{1}_{z_{1}}\times\cdots\times\mathbb{P}^{1}_{z_{n}},\sum_{i}((z_{i}=0)+(z_{i}=\infty))).$
  4. (4)
    The pair $(X,D_{X})$ is toric.
A result analogous to Theorem 1.1 or Theorem 1.2 was obtained by Cadorel, Deng, and Yamanoi [Deng2024HyperFund] for Campana’s special manifolds. Recall that projective manifolds with nef anticanonical bundle are special (cf. [Campana2004Orbifolds, Example 2.3]). Moreover, Shokurov’s conjecture [Shokurov2000CompleSurf] only assumes the nefness of $(K_{X}+D_{X})$. It is therefore natural to expect that the main Theorems of this note remain valid for log canonical pairs with $(K_{X}+D_{X})$ nef.
The note is organized as follows. In Section 2, we recall the relevant definitions and develop the necessary technical tools. In Section 3, we prove the main theorems.
Acknowledgments
The author thanks professors Zhan Li, Lei Zhang, Junchao Shentu, and Osamu Fujino for helpful discussions or comments. The author thanks professor Ya Deng for drawing his attention to related results in joint work with Benoit Cadorel and Katsutoshi Yamanoi concerning hyperbolicity and fundamental groups for quasi-projective varieties. The author is supported by USTC.
2. Preparations
We work over the field of complex numbers. For the singularities of pairs and the minimal model program, we refer to [KM98, Kollar2013sing].
A dominant morphism between normal varieties with irreducible general fibers is called a fibration. Let $f:X\to Y$ be a projective fibration between normal varieties. Let $D$ be an effective $\mathbb{Q}$-divisor on $X$. We can write $D=D^{h}+D^{v}$, where $D^{v}$ does not dominate $Y$, and each component of $D^{h}$ dominates $Y$. $D^{v}$ (resp. $D^{h}$) is called the vertical (resp. horizontal) part of $D$. $D$ is called vertical (resp. horizontal) if $D=D^{v}$ (resp. $D=D^{h}$). Assume that $D$ is vertical, then $D$ is called exceptional over $Y$ if $\operatorname{codim}f(D)\geq 2$, and degenerate over $Y$ if for any prime divisor $P\subset Y$, there is a prime divisor $Q\subset X$ such that $Q$ dominates $P$ but $Q\not\subset D$.
For a birational map $\pi:X\dashrightarrow Y$ between normal varieties, we set $\operatorname{Ex}(\pi)=\sum P$, where the sum is taken over all prime divisors on $X$ which are exceptional over $Y$. A birational map $\pi:X\dashrightarrow Y$ is called a birational contraction if $\operatorname{Ex}(\pi^{-1})=0$.
For a projective variety $X$, we set $q(X)=\dim H^{1}(X,O_{X})$, and for a projective lc pair $(X,D_{X})$, we set $q(X,D_{X})=\dim H^{0}(X^{0},\Omega_{X^{0}}^{1}(\log\lfloor D_{X}|_{X^{0}}\rfloor))$, where $X^{0}$ is the log smooth locus of $(X,D_{X})$.
2.1. Differential forms on lc pairs
Theorem 2.1 ([greb2011differential]).
Let $(X,D_{X})$ be an lc pair. Let $X^{0}$ be its log smooth locus and let $\operatorname{nklt}(X,D_{X})$ be its non-klt locus. Let $\pi:Y\to X$ be a log resolution of $(X,D_{X})$ and let $D_{Y}$ be the codimension one part of $\pi^{-1}\operatorname{nklt}(X,D_{X})$. Then $\pi_{*}\Omega_{Y}^{k}(\log D_{Y})$ is reflexive and $H^{0}(Y,\Omega_{Y}^{k}(\log D_{Y}))\cong H^{0}(X^{0},\Omega_{X^{0}}^{k}(\log\lfloor D_{X}|_{X^{0}}\rfloor))$ for all $k$.
We will use the following Lemma repeatedly.
Lemma 2.2.
Let $(X,D_{X})$ be a projective lc pair. Let $\pi:Y\to X$ be a log resolution of $(X,D_{X})$ and let $D_{Y}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. Then the following holds.
  1. (1)
    $q(X,D_{X})=q(Y,D_{Y})$.
  2. (2)
    $\kappa(K_{X}+D_{X})=\kappa(K_{Y}+D_{Y})$ and $\nu(K_{X}+D_{X})=\nu(K_{Y}+D_{Y})$, where $\kappa$ and $\nu$ are Kodaira dimension and numerical Kodaira dimension respectively.
  3. (3)
    If $(X,D_{X})$ is dlt, then $q(X)=q(Y)$.
Proof.
Since $(X,D_{X})$ is lc, there is an effective exceptional divisor $E$ on $Y$ such that $K_{Y}+D_{Y}=\pi^{*}(K_{X}+D_{X})+E$, so we have (2). If $(X,D_{X})$ is dlt, then $X$ has rational singularities by [KM98, Theorem 5.22]. Hence we have (3).
Let $D_{Y}^{\prime}$ be the codimension one part of $\pi^{-1}\operatorname{nklt}(X,D_{X})$. Note that $D_{Y}^{\prime}\leq\lfloor D_{Y}\rfloor$. To show (1), it suffices to show that the injective homomorphism
$\iota:H^{0}(Y,\Omega_{Y}^{1}(\log D_{Y}^{\prime}))\to H^{0}(Y,\Omega_{Y}^{1}(\log\lfloor D_{Y}\rfloor))$
is an isomorphism. This isomorphism is a special case of [greb2011differential, Proposition 19.1]. Here we give this isomorphism a proof using Hacon’s $\mathbb{Q}$-factorial dlt modification [Hacon2014ACC, Proposition 3.3.1].
Step 1. Assume that $(X,D_{X})$ is $\mathbb{Q}$-factorial dlt. Then $\operatorname{nklt}(X,D_{X})=\lfloor D_{X}\rfloor$ and $D_{Y}^{\prime}=q^{-1}\lfloor D_{X}\rfloor$. Write $D_{Y}^{\prime}=\sum_{i}P_{i}$ and $\lfloor D_{Y}\rfloor-D_{Y}^{\prime}=\sum_{j}E_{j}$ as sums of prime divisors. By the exact sequence
one checks that $\iota$ is an isomorphism iff $\sum_{i}a_{i}c_{1}(O_{Y}(P_{i}))+\sum_{j}b_{j}c_{1}(O_{Y}(E_{j}))=0$ implies that all $b_{j}$’s are zero (see Lemma 3.7). This is indeed the case since all $E_{j}$’s are exceptional and $X$ is $\mathbb{Q}$-factorial.
Step 2. Assume that $(X,D_{X})$ is log smooth. By Iitaka [iitaka1977logarithmicKodaira, Proposition 1], we have $q(X,\lfloor D_{X}\rfloor)=q(Y,\pi^{-1}\lfloor D_{X}\rfloor)$. Hence we have $q(X,\lfloor D_{X}\rfloor)=q(Y,\lfloor D_{Y}\rfloor)$ by Step 1, i.e., (1) holds for log smooth pairs.
Step 3. In general, consider the following diagram.
@@ 2.1. Differential forms on lc pairs
where
  • $(Z,D_{Z})$ is a good minimal model of $(Y,D_{Y})$ over $X$, hence a Hacon’s $\mathbb{Q}$-factorial dlt modification of $(X,D_{X})$ [hacon2007shokurov, Proposition 3.3.1], and
  • $Y\xleftarrow{p}W\xrightarrow{q}Z$ is a common log resolution.
Since $K_{Z}+D_{Z}=\mu^{*}(K_{X}+D_{X})$, we have $\mu(\lfloor D_{Z}\rfloor)=\operatorname{nklt}$. Hence $\lfloor D_{Z}\rfloor\subset\mu^{-1}\operatorname{nklt}$ and $q^{-1}\lfloor D_{Z}\rfloor\subset(\mu\circ q)^{-1}\operatorname{nklt}$. Let $D_{W}^{\prime}$ be the codimension one part of $(\mu\circ q)^{-1}\operatorname{nklt}$, then $q^{-1}\lfloor D_{Z}\rfloor\leq D_{W}^{\prime}$. Let $D_{W}=q^{-1}_{*}D_{Z}+\operatorname{Ex}(q)$. Then $D_{W}=(\mu\circ q)^{-1}_{*}D_{X}+\operatorname{Ex}(\mu\circ q)$. By $q^{-1}\lfloor D_{Z}\rfloor\leq D_{W}^{\prime}\leq\lfloor D_{W}\rfloor$ we have $q(W,q^{-1}\lfloor D_{Z}\rfloor)\leq q(W,D_{W}^{\prime})\leq q(W,\lfloor D_{W}\rfloor)$. By Step 1, we have $q(W,q^{-1}\lfloor D_{Z}\rfloor)=q(W,\lfloor D_{W}\rfloor)$. Hence
(2.1.1) $q(W,D_{W}^{\prime})=q(W,\lfloor D_{W}\rfloor).$
By construction, we have $p^{-1}D_{Y}^{\prime}\subset D_{W}^{\prime}\subset p^{-1}_{*}D_{Y}^{\prime}+\operatorname{Ex}(p)$. Hence $q(W,p^{-1}D_{Y}^{\prime})\leq q(W,D_{W}^{\prime})\leq q(W,p^{-1}_{*}D_{Y}^{\prime}+\operatorname{Ex}(p))$. By step 1, we have $q(W,p^{-1}D_{Y}^{\prime})=q(W,p^{-1}_{*}D_{Y}^{\prime}+\operatorname{Ex}(p))$. Hence $q(W,D_{W}^{\prime})=q(W,p^{-1}D_{Y}^{\prime})$. By Iitaka [iitaka1977logarithmicKodaira, Proposition 1], we have
(2.1.2) $q(W,D_{W}^{\prime})=q(W,p^{-1}D_{Y}^{\prime})=q(Y,D_{Y}^{\prime})$
By Step 2, we have
(2.1.3) $q(W,\lfloor D_{W}\rfloor)=q(Y,\lfloor D_{Y}\rfloor)$
Hence $q(Y,D_{Y}^{\prime})=q(Y,\lfloor D_{Y}\rfloor)$ by the three equations 2.1.1, 2.1.2 and 2.1.3. ∎
Lemma 2.3.
Let $(X,D_{X})$ and $(Y,D_{Y})$ be two projective lc pairs. Assume that there is a birational contraction $\pi:X\dashrightarrow Y$ such that $D_{X}=\pi^{-1}_{*}D_{Y}+\operatorname{Ex}(\pi)$ then $q(X,D_{X})=q(Y,D_{Y})$.
@@
Let $X\xleftarrow{p}W\xrightarrow{q}Y$ be a common log resolution. Then $q^{-1}_{*}D_{Y}+\operatorname{Ex}(q)=p^{-1}_{*}D_{X}+\operatorname{Ex}(p)$. Hence, the Lemma follows from Lemma 2.2. ∎
2.2. Quasi-Albanese morphisms for smooth varieties
Theorem 2.4 ([iitaka1976logarithmic], see also [fujino2024quasi]).